OpenStudy (biue):

John’s piggy bank contained 7 dimes and 3 quarters. He pulled out 1 coin without looking. Without replacing the first coin, John then pulled out a second coin. What is the probability that both coins were dimes? Express your answer in simplest form.

OpenStudy (biue):

its supposed to be a fraction as the answer, but I'm having trouble finding it out.

rebeccaxhawaii (rebeccaxhawaii):

okay so 7:3

OpenStudy (biue):

?

OpenStudy (biue):

@rebeccaxhawaii

OpenStudy (biue):

7/15/?!??!

rebeccaxhawaii (rebeccaxhawaii):

the chances of pulling 2 dimes is 2 out of 12 because there is 12 coins in all

rebeccaxhawaii (rebeccaxhawaii):

Please click "best response" for the person that helped you the most, this just shows the appreciation for the time they took to assist. If you want further help with anymore questions, please fan me so in the long run you can see when I'm on. I'll be sure to take a look at your question to help :)

OpenStudy (sydniek17):

P(A)=dime on first try P(B)=dime on second try 7 dimes+3 quarters=10 total coins P(A)=7/10 Now there are only 6 dimes and 3 quarters, equaling 9 total coins. P(B)=6/9=2/3 Since this is an "and" problem, we will multiply the two probabilities: P(A)*P(B) (7/10)*(2/3)=14/30 14/30 simplifies to your final answer of 7/15.